a firm wishes to maintain an internal growth rate of 6.75 percent and a dividend payout ratio of 31 percent. the current profit margin is 5.3 percent and the firm uses no external financing sources

what must total asset turnover be
Answer #1

Retention ratio = 1 - 0.31 = 0.69

Internal growth rate = 0.0675 = (ROA × 0.69)/[1 - (ROA × 0.69)]

ROA = 0.09164

Return on assets = 0.09164 = 0.053 × TAT

Total asset turnover = 1.73 times